# B-VLLM: A Vision Large Language Model with Balanced Spatio-Temporal Tokens

### v1

@@ Abstract
Recently, Vision Large Language Models (VLLMs) integrated with vision encoders have shown promising performance in vision understanding. The key of VLLMs is to encode visual content into sequences of visual tokens, enabling VLLMs to simultaneously process both visual and textual content. However, understanding videos, especially long videos, remain a challenge to VLLMs as the number of visual tokens grows rapidly when encoding videos, resulting in the risk of exceeding the context window of VLLMs and introducing heavy computation burden. To restrict the number of visual tokens, existing VLLMs either: (1) uniformly downsample videos into a fixed number of frames or (2) reducing the number of visual tokens encoded from each frame. We argue the former solution neglects the rich temporal cue in videos and the later overlooks the spatial details in each frame. In this work, we present Balanced-VLLM (B-VLLM): a novel VLLM framework that aims to effectively leverage task relevant spatio-temporal cues while restricting the number of visual tokens under the VLLM context window length. At the core of our method, we devise a text-conditioned adaptive frame selection module to identify frames relevant to the visual understanding task. The selected frames are then de-duplicated using a temporal frame token merging technique. The visual tokens of the selected frames are processed through a spatial token sampling module and an optional spatial token merging strategy to achieve precise control over the token count. Experimental results show that B-VLLM is effective in balancing the number of frames and visual tokens in video understanding, yielding superior performance on various video understanding benchmarks. Our code is available at https://github.com/zhuqiangLu/B-VLLM.
@@ 1 Introduction
Figure 1: Illustration of the limitation of existing VLLMs.
@@
Large Language Models (LLMs) have demonstrated remarkable performance and generalizability in natural language understanding [3, 44, 40, 8]. Recently, there have been several notable efforts to extend LLMs into Vision Large Language Models (VLLMs) for visual understanding [1, 27, 48]. These efforts primarily aim to adapt LLMs’ contextual reasoning capabilities to multimodal settings like video captioning and visual question answering, enabling them to generate accurate textual responses based on visual and textual context provided by users. To bridge the modality gap between textual and visual contexts, existing VLLMs typically use pretrained Vision Transformers (ViTs) [7] to encode images into sequences of visual tokens. These tokens are subsequently integrated with text token sequences as input to the VLLMs [27, 26]. This approach compresses spatial contexts from images into 1D sequences and has demonstrated its effectiveness across various computer vision tasks [11].
@@
However, these VLLMs are primarily designed for single-image scenarios. As illustrated in Figure 1(a), extending them to video data often compromises performance, as the number of visual tokens exceeds the VLLM’s context window limit. To address this issue, methods like LLaMA-VID [24], MovieChat [39] and LongVLM [42] reduce the number of visual tokens to encode each frame. Alternatively, methods such as VideoLLaMA2 [6], VideoLLaMA [45] and VideoChat2 [21] sample a fixed number of frames uniformly from a video to prevent an overload of visual tokens. While being effective in controlling the number of visual tokens for video processing, these methods struggle to balance spatio-temporal cues. As depicted in Figure 1(b), VLLMs using uniform frame sampling tend to underperform on temporally relevant tasks. As shown in Figure 1(c), methods that reduce the number of visual tokens per frame often perform poorly on spatially relevant tasks. We argue that this is due to a spatio-temporal token imbalance: reducing frame-level visual tokens causes temporal cues to dominate, whereas uniform frame sampling leads to spatial cues to be overshadowed.
@@
To address these issue, we propose Balanced-VLLM (B-VLLM), a VLLM framework designed to balance spatio-temporal cues in video data conditioned by task related text prompts. B-VLLM processes each frame with both coarse-grained and fine-grained visual tokens. A learnable frame selection module leverages coarse-grained tokens to identify the frames most relevant to the video understanding task, guided by the text prompt. The fine-grained visual tokens of these selected frames are then processed by a spatial visual token sampling module to retrieve those most task-relevant ones. To further control the final number of generated visual tokens, we introduce an iterative spatial token merging strategy that averages similar visual tokens until the desired number of tokens is reached. Finally, the resulting visual tokens are mapped into the LLM feature space and integrated with the sequence of textual tokens encoded from the text prompt before being fed into the LLM. Notably, our proposed method achieves SOTA performance on various benchmarks. Especially, our method introduces up to 10% performance improvement to the backbone VLLMs in video understanding.
In summary, the main contributions of this paper are as follows:
@@
  • We undertake comprehensive experiments to show that B-VLLM is effective in different VLLM architectures for video understanding regime.
@@
2 Related Work
2.1 Vision Large Language Model
Vision Large Language Models (VLLMs) extend existing Large Language Models(LLMs) to the domain of visual understanding through visual-instruction tuning [27]. Visual content is encoded as sequences of visual tokens using a pretrained CLIP [35] vision encoder. These visual tokens are then projected into the LLM hyperspace through a trainable MLP layer [27, 24, 6, 25]. However, in multi-image or video scenarios, the number of vision tokens often exceeds the VLLM’s context window limit, leading to performance degrading and high computation costs. To address this, Open-Flamingo[1], MovieChat [39] and BLIP2  [20] reduce the number of vision tokens extracted from each frame by using transformer-based token selection networks, such as Qformer [20] and Perceptive Resampler [1].
However, the number of vision tokens is proportional to the number of input images or frames, creating a risk of an overwhelming number of tokens when processing long videos. LLaMA-VID [24] addresses this by encoding each frame into only two visual tokens to enable hour-long video processing; however, this extreme compression rate limits LLaMA-VID’s ability to capture finer spatial details in each frame. Another common approach to controlling the number of visual tokens is to uniformly sample frames from a video [27, 25, 26, 41, 6]. However, uniformly frame sampling may overlook important frames, especially when keyframes of interests are not evenly distributed throughout the video. We argue that frames most relevant to a given visual understanding task often concentrate in specific segments of the video. Uniformly sampling may skip these critical cues, leading to suboptimal performance.
Compared to existing studies, we propose a vision LLM framework empowered by an adaptive, text-conditioned frame selection mechanism that dynamically locates and selects the most relevant frames for the given vision task.
2.2 Token Merging Techniques
Token merging [5] is a technique that combines similar tokens to reduce computational costs with minimum impact on performance in vision tasks. Concurrently, SD-TOME [4] significantly reduces the computation load of text-to-image generation models through token merging. VID-TOME [22] demonstrates that token merging is also effective for processing visual tokens in videos.
In LLMs [17], token merging significantly improves the inference speed by reducing the number of textual tokens. In VLLMs, TOME-LLaVA [37] merges visual tokens to lower computational cost. More recently, Chat-Uni-VI [18] integrates token merging into video LLM pipeline to reduce the number of visual tokens for long videos. LongVLM [42] divides long videos into segments and applies token merging within each segment’s visual tokens. Despite these studies integrating token merging into VLLMs, they often lacks flexibility in controlling the number of tokens. Our study for the first time proposes an iterative token merging strategy to achieve finer control over the number of tokens.
@@ 3.1 Architecture Overview
First, our text-conditioned frame selection module utilizes the coarse semantics captured by $L$ [CLS] tokens in the video and the given text context $T$ to select the $L^{*}$ most informative frames. The selected frames with their visual tokens can be denoted as: $V^{*}=\{V^{*}_{1},..,V^{*}_{L^{*}}\}\subseteq V$. To further lower the total number of visual tokens, we utilize a frame-wise token selection module to sample the $Z$ most informative visual tokens from each $V^{*}_{l^{*}}\in V^{*}$, conditioning to $T$. An optional spatial token merging module is proposed to combine the selected visual tokens that bear high similarity for a further control in the number of tokens. The final selected visual tokens are projected into the LLM feature space and combined with the text context $T$ to form as the input of LLM for downstream tasks, such as video captioning and visual question answering.
@@ 3.2 Text-Conditioned Adaptive Frame Selection
Using only the frames most relevant to the given text context is essential for reducing the number of visual tokens. However, identifying these frames can be computationally expensive. To address this, we use the [CLS] token $V_{l,[CLS]}$ for each frame instead of including all other preliminary frame-level visual tokens $V_{l,1},...V_{l,N}$. This design choice balances computational efficiency with the effectiveness of locating relevant frames. We adapt a QFormer [20] as our frame selection network due to its lightweight design and multimodal reasoning capability.
In detail, the QFormer creates $L^{*}$ queries $Q=\{q_{1},...,q_{L^{*}}\}$ via jointly encoding the [CLS] tokens $V_{[CLS]}=\{V_{1,[CLS]},...V_{L,[CLS]}\}$ and the text context $T$. These queries are for locating the $L^{*}$ most informative frames from $L^{*}$ perspectives and their corresponding visual tokens. Mathematically, we have:
@@
Notably, as $\tau$ approaches 0, the Gumbel-Softamx distribution converges to a categorical distribution, meaning that $S_{\tau=0}$ is a matrix where each row is a one-hot vector. Leveraging the property of the Gumbel-Softmax, we construct $\bar{V}$ to enable discretely sampling from $V$. Hence, conceptually, we have $V^{*}_{l^{*}}\in V^{*}$. In particular, if $s_{\tau=0,l^{*},l}=1$, this indicates that ${V}_{l^{*}}^{*}=V_{l}$, meaning $V_{l}$ is selected for the $l^{*}$-$th$ perspective. In this work, we typically set $\tau$ to 0.5 to ensure the a close approximation of one-hot vectors. We argue that Softmax is unsuitable in this context as $S_{\tau}$ is calculated based on $V_{[CLS]}$ where each $V_{[CLS],l}$ reflects coarse semantic and may overlook fine-grained spatial details. As a result, Softmax may smooth out rich spatial visual cues especially when aggregating multiple $V_{l}$s.
@@ 3.3 Temporal Frame Token Merging
The text-conditioned adaptive frames selection retrieves the $L^{*}$ frames with their preliminary visual tokens that are most relevant to the context $T$. However, duplicated contents in the selected frames may occur, for example, when the number of frames $L$ is less than $L^{*}$, leading to duplicated frames. To address this, we propose a temporal visual token merging mechanism to combine similar or duplicated frames selected.
Starting with duplication detection, we use $S_{\tau}$ to identify duplicates in the selected frames. As each perspective (i.e., row-vector) in $S_{\tau}$ approximates a one-hot vector distribution, thereby we can measure the row-wise similarity to locate the potential duplications. With a cosine similarity measure, two row vectors are marked as duplicates if their similarity exceeds a pre-set threshold $\gamma$.
We conduct this procedure in an iterative manner. For $V^{*}_{\alpha}\in V^{*}$, we identify and update the frames iteratively as a set:
@@ 3.4 Spatial Visual Token Sampling
Although the maximum cardinality of $V^{*}$ is limited, each $V^{*}_{i}$ contains $M$ visual tokens except the [CLS] token, resulting in a total of up to $L^{*}M$ visual tokens. Existing studies [27, 46] typically set $M$ to 256 or 576. When 32 frames are selected, this result in 8,192 or 18,432 visual tokens, respectively, increasing the risk of exceeding the LLM context window limit and imposing a heavy computation load. To reduce the number of visual tokens per frame, we use a spatial QFormer to sample $R$ visual tokens that are most relevant to the given context $T$, with $R\ll M$. In detail, for each $V^{*}_{l^{*}}$, the spatial Qformer generates $R$ learnable spatial visual tokens: $V^{**}_{l^{*}}=\{V^{**}_{l^{*},1},...,V^{**}_{l^{*},R}\}$ with the inputs $V^{*}_{l^{*}}=\{V^{*}_{l^{*},1},...,V^{*}_{l^{*},M}\}$ and the text context $T$. Finally, each video contains a total of $MR$ visual tokens.
Notably, the spatial Qformer generates a fixed number of tokens, which limits flexibility when finer control over the token count is needed. To address this, we propose an optional progressive spatial visual token merging strategy for the scenarios where a maximum visual token count $\theta$ applies. Simply put, if the size of $V^{**}_{l^{*}}$ exceeds the maximum token limit, we halve $V^{**}_{l*}$ by merging the most similar tokens through Bipartite Merging [5]. This process is repeated until the number of tokens is satisfied.
@@ 3.5 Integration with Backbone LLMs
As the backbone LLM is a unimodal model that only responds to textual input, we bridge the modality gap between textual and the visual tokens via a learnable MLP model. Specifically, we utilize a one-layer MLP module without an activation function to linearly map the visual tokens from our proposed modules to the LLM’s feature space. Finally, the projected visual tokens are concatenated with the textual token sequence before being fed to the backbone LLM. Notably, our proposed method is flexible in either integrating with existing VLLM or in being a VLLM on its own. In particular, it is compatible with most of the existing VLLMs architectures, such as LLaMA-VID [24] and VideoLLaMA2 [6]. Here, we denote our method as B-VLLM when there is no integration with other VLLM. In such case, we utilize Qwen2 [44] as backbone LLM for B-VLLM.
4 Implementation
Baselines. For a comprehensive evaluation, we adopt two baselines - LLaMA-VID [24] and VideoLLaMA2 [6]. LLaMA-VID is a representative VLLM that achieves video understanding with a high spatial compression rate. We follow the official LLaMA-VID setting that each frame is encoded with two visual tokens. In contrast, VideoLLaMA2 uniformly samples frames from videos. Following the official setting, 8 frames are uniformly sampled from each video and visual tokens are subsequently downsampled by RegStage [36] and a series of 3D convolution blocks, in total 576 visual tokens are obtained.
Datasets. We report the performances of B-VLLM exclusively trained on two open-source datasets: LLaMA-VID-Dataset [24] and Valley [31] for a fair comparison with the two baselines. Please refer to supplementary material for the details of datasets.
Training. We adopt a two-stage training strategy [27, 24, 6], dividing training into pretraining for modality alignment and fine-tuning for instruction tuning. During pretraining, the learning rate is set to 1e-3 with a linear learning rate schedule, and the total batch size is set to 256. Only the frame selection module and the MLP projection layer, which maps the visual token features to the LLM features, are updated. In the fine-tuning stage, the entire framework is unfrozen except for the preliminary visual encoder. Additionally, LoRA [15] is enabled with the rank set to 128 and $\alpha$ set to 256. The total batch size was set to 128, and the learning rate is 1e-4 with cosine scheduling. Supervision is provided by minimizing cross-entropy for masked text tokens, and the optimizer for both stages is AdamW [29], with DeepSpeed ZeRO2 [2] enabled for memory efficiency. Training is conducted with 16 Nvidia L40S GPUs. Following previous works [21, 32, 6], videos are downsampled to 1 fps in both training and evaluation.
5 Experimental Result & Discussion
@@
$\Delta$ +1.0 +2.4 +1.8 +1.9 -1.7 +1.7 +3.8 +1.9/+0.1 +3.3/0.2 +0.9/+0.0
B-VLLM 1fps 49.2 50.8 13.8 51.6 59.1 52.0 44.4 72.3/4.0 58.2/3.5 45.3/3.4

### v2

@@ Abstract
Recently, Vision Large Language Models (VLLMs) with integrated vision encoders have shown promising performance in vision understanding. They encode visual content into sequences of visual tokens, enabling joint processing of visual and textual data. However, understanding videos, especially long videos, remains a challenge as the rapid growth of visual tokens during video encoding risks exceeding VLLMs’ context window length and significantly escalates computational cost. To restrict the number of visual tokens, existing VLLMs either: (1) uniformly downsample videos into a fixed number of frames or (2) reducing the number of visual tokens encoded from each frame. We argue that the former neglects temporal dynamics in videos, while the latter fails to preserve spatial details within individual frame. In this work, we propose Balanced-VLLM (B-VLLM), a novel VLLM framework designed to model task relevant spatio-temporal cues, while restricting the number of visual tokens within the VLLM’s context window length. Central to our framework is a text-conditioned adaptive frame selection module that dynamically identifies task-relevant frames, which are further de-duplicated with a temporal frame token merging strategy. The visual tokens of these frames then undergo spatial token sampling and an optional spatial token merging strategy for granular control against the token budget. Experiments demonstrate the effectiveness of B-VLLM in balancing the number of frames and visual tokens, moreover, our proposed method introduces 10% performance gain on MVBench. Code is available at https://github.com/zhuqiangLu/B-VLLM.git.
22footnotetext: Corresponding authors.
@@ 1 Introduction
Figure 1: Illustration of the limitations of existing VLLMs.
@@
Large Language Models (LLMs) have demonstrated remarkable capability and generalizability in natural language understanding [4, 56, 51, 11]. Recently, there have been several notable efforts to extend LLMs into Vision Large Language Models (VLLMs) for visual understanding [2, 33, 62]. These efforts primarily aim to adapt LLMs’ contextual reasoning capabilities to multimodal settings like video captioning and visual question answering, enabling them to generate accurate textual responses based on visual and textual context provided by users. Unlike the extension of vision models to downstream tasks [19, 57, 37, 61, 8], bridging the modality gap between textual and visual contexts poses significant challenges [33]. Existing VLLMs typically use pretrained Vision Transformers (ViTs) [10] to encode images into sequences of visual tokens. These tokens are subsequently integrated with text token sequences as input to the VLLMs [33, 32]. This approach compresses spatial contexts from images into 1D sequences and has demonstrated its effectiveness across various computer vision tasks [14]. However, these VLLMs are primarily designed for single-image scenarios. As illustrated in Figure 1(a), extending them to video data often compromises performance, as the number of visual tokens exceeds the length limit of the VLLM’s context window.
@@
To address this issue, methods like LLaMA-VID [30], MovieChat [50] and LongVLM [54] reduce the number of visual tokens to encode each frame. Alternatively, methods such as VideoLLaMA2 [9], VideoLLaMA [58] and VideoChat2 [27] sample a fixed number of frames uniformly from a video to prevent an overload of visual tokens. While being effective in controlling the number of visual tokens for video processing, these methods struggle to balance spatio-temporal cues. As depicted in Figure 1(b), VLLMs using uniform frame sampling tend to underperform on temporally relevant tasks. As shown in Figure 1(c), methods that reduce the number of visual tokens per frame often perform poorly on spatially relevant tasks.
@@
We argue that this is due to a spatio-temporal token imbalance: reducing frame-level visual tokens causes temporal cues to dominate, whereas uniform frame sampling leads to spatial cues to be overshadowed. To address these issues, we propose Balanced-VLLM (B-VLLM), a novel VLLM framework to balance spatio-temporal cues in video data conditioned by task-related text prompts. B-VLLM processes each frame with both coarse-grained and fine-grained visual tokens. A learnable frame selection module leverages coarse-grained tokens to identify the frames most relevant to the video understanding task, guided by the text prompt. The fine-grained visual tokens of these selected frames are then processed by a spatial visual token sampling module to retrieve those most task-relevant ones. For granular control over the final number of generated visual tokens, we introduce an iterative spatial token merging strategy that averages similar visual tokens until the expected number of tokens is reached. Finally, the resulting visual tokens are mapped into the LLM feature space and integrated with the sequence of textual tokens encoded from the text prompt before being fed into the LLM. Our proposed method achieves SOTA performance on various benchmarks. Notably, our method introduces up to 10% performance improvement on MVBench [27]
The main contributions are summarized as follows:
@@
  • We undertake comprehensive experiments to show that B-VLLM is effective in different VLLM architectures for video understanding.
@@
2 Related Work
2.1 Vision Large Language Models
Vision Large Language Models (VLLMs) extend existing Large Language Models (LLMs) to the domain of visual understanding through visual-instruction tuning [33]. Visual content is encoded as sequences of visual tokens using a pretrained CLIP [44] vision encoder. These visual tokens are then projected into the LLM hyperspace through a trainable MLP layer [33, 30, 9, 31]. However, in multi-image or video scenarios, the number of vision tokens often exceeds the VLLM’s context window limit, leading to performance degrading and high computation costs. To address this, Open-Flamingo[2], MovieChat [50] and BLIP2  [26] reduce the number of vision tokens extracted from each frame by using transformer-based token selection networks, such as Q-Former [26] and Perceptive Resampler [2]. However, the number of vision tokens is proportional to the number of input images or frames, creating the risk of an overwhelming number of tokens when processing long videos. LLaMA-VID [30] addresses this by encoding each frame into only two visual tokens to enable hour-long video processing; however, this extreme compression rate limits LLaMA-VID’s ability to capture finer spatial details in each frame. Another common approach to controlling the number of visual tokens is to uniformly sample frames from a video [33, 31, 32, 53, 9]. However, uniformly frame sampling may overlook important frames, especially when keyframes of interests are not evenly distributed throughout the video. Compared to existing studies, we propose a vision LLM framework empowered by an adaptive, text-conditioned frame selection mechanism that dynamically selects the most relevant frames for the given vision task.
2.2 Token Merging Techniques
Token merging [7] is a technique that combines similar tokens to reduce computational costs with minimum impact on performance in vision tasks. Concurrently, SD-TOME [6] significantly reduces the computation load of text-to-image generation models through token merging. VID-TOME [28] demonstrates that token merging is also effective for processing visual tokens in videos. In LLMs [21], token merging significantly improves the inference speed by reducing the number of textual tokens. In VLLMs, TOME-LLaVA [46] merges visual tokens to lower computational cost. More recently, Chat-Uni-VI [22] integrates token merging into video LLM pipeline to reduce the number of visual tokens for long videos. LongVLM [54] divides long videos into segments and applies token merging within each segment’s visual tokens. Despite these studies integrating token merging into VLLMs, they often lack flexibility in controlling the number of tokens. Our study for the first time proposes an iterative token merging strategy to achieve finer control over the number of tokens.
@@ 3.1 Architecture Overview
First, our text-conditioned frame selection module utilizes the coarse semantics captured by $L$ [CLS] tokens in the video and the given text context $T$ to select the $L^{*}$ most informative frames. The selected frames with their visual tokens can be denoted as: $V^{*}=\{V^{*}_{1},..,V^{*}_{L^{*}}\}\subseteq V$. To further lower the total number of visual tokens, we utilize a frame-wise token selection module to sample the $Z$ most informative visual tokens from each $V^{*}_{l^{*}}\in V^{*}$, conditioning on $T$. An optional spatial token merging module is proposed to combine the selected visual tokens that bear high similarity for a further control in the number of tokens. The final selected visual tokens are projected into the LLM feature space and combined with the text context $T$ to form as the input of LLM for downstream tasks, such as video captioning and visual question answering.
@@ 3.2 Text-Conditioned Adaptive Frame Selection
Using only the frames most relevant to the given text context is essential for reducing the number of visual tokens. However, identifying these frames can be computationally expensive. To address this, we use the [CLS] token $V_{l,[CLS]}$ for each frame instead of including all other preliminary frame-level visual tokens $V_{l,1},...V_{l,N}$. This design choice balances computational efficiency with the effectiveness of locating relevant frames. We adapt a Q-Former [26] as our frame selection network due to its lightweight design and multimodal reasoning capability.
In detail, the Q-Former creates $L^{*}$ queries $Q=\{q_{1},...,q_{L^{*}}\}$ via jointly encoding the [CLS] tokens $V_{[CLS]}=\{V_{1,[CLS]},...V_{L,[CLS]}\}$ and the text context $T$. These queries are for locating the $L^{*}$ most informative frames from $L^{*}$ perspectives and their corresponding visual tokens. Mathematically, we have:
@@
Notably, as $\tau$ approaches 0, the Gumbel-Softamx distribution converges to a categorical distribution, meaning that $S_{\tau=0}$ is a matrix where each row is a one-hot vector. Leveraging the property of the Gumbel-Softmax, we construct $\bar{V}$ to enable discretely sampling from $V$. Hence, conceptually, we have $V^{*}_{l^{*}}\in V^{*}$. In particular, if $s_{\tau=0,l^{*},l}=1$, this indicates that ${V}_{l^{*}}^{*}=V_{l}$, meaning $V_{l}$ is selected for the $l^{*}$-thth perspective. In this work, we typically set $\tau$ to 0.1 to ensure the a close approximation of one-hot vectors. Lastly, we restore the temporal order of $V^{*}$ based on $S_{\tau}$ as Gumbel-Softmax may break the order of selected frames. We argue that Softmax is unsuitable in this context as $S_{\tau}$ is calculated based on $V_{[CLS]}$ where each $V_{[CLS],l}$ reflects coarse semantic and may overlook fine-grained spatial details. As a result, Softmax may smooth out rich spatial visual cues especially when aggregating multiple $V_{l}$s.
3.2.1 Discussion: Frame Selection with [CLS] Token.
The role of [CLS] tokens, which carry high-level semantic information of corresponding images [52, 63], has been widely studied, and our ablation studies in Section 5.3 validate their effectiveness in frame selection.
@@ 3.3 Temporal Frame Token Merging
The text-conditioned adaptive frames selection retrieves the $L^{*}$ frames with their preliminary visual tokens that are most relevant to the context $T$. However, duplications in the selected frames may occur, for example, when the number of frames $L$ is less than $L^{*}$, leading to duplications. To address this, we propose a temporal visual token merging mechanism to combine similar or duplicated frames selected.
Starting with duplication detection, we use $S_{\tau}$ to identify duplicates in the selected frames. As each perspective (i.e., row-vector) in $S_{\tau}$ approximates a one-hot vector distribution, thereby we can measure the row-wise similarity to locate the potential duplications. With a cosine similarity measure, two row vectors are marked as duplicates if their similarity exceeds a pre-set threshold $\gamma$. We conduct this procedure in an iterative manner. For $V^{*}_{\alpha}\in V^{*}$, we identify and update the frames iteratively as a set:
@@ 3.4 Spatial Visual Token Sampling
Although the maximum cardinality of $V^{*}$ is limited, each $V^{*}_{i}$ contains $M$ visual tokens except the [CLS] token, resulting in a total of up to $L^{*}M$ visual tokens. Existing studies [33, 59] typically set $M$ to 256 or 576. When 32 frames are selected, this result in 8,192 or 18,432 visual tokens, respectively, increasing the risk of exceeding the LLM context window limit and imposing a heavy computation load. To reduce the number of visual tokens per frame, we use a spatial Q-Former to sample $R$ visual tokens that are most relevant to the given context $T$, with $R\ll M$. In detail, for each $V^{*}_{l^{*}}$, the spatial Q-Former generates $R$ learnable spatial visual tokens: $V^{**}_{l^{*}}=\{V^{**}_{l^{*},1},...,V^{**}_{l^{*},R}\}$ with the inputs $V^{*}_{l^{*}}=\{V^{*}_{l^{*},1},...,V^{*}_{l^{*},M}\}$ and the text context $T$. Finally, each video contains a total of $MR$ visual tokens.
Notably, the spatial Q-Former generates a fixed number of tokens, which limits flexibility when finer control over the token count is needed. To address this, we propose an optional progressive spatial visual token merging strategy for the scenarios where a maximum visual token count $\theta$ applies. Simply put, if the size of $V^{**}_{l^{*}}$ exceeds the maximum token limit, we halve $V^{**}_{l*}$ by merging the most similar tokens through Bipartite Merging [7]. This process is repeated until the number of tokens is satisfied.
@@ 3.5 Integration with Backbone LLMs
As the backbone LLM is a unimodal model that only responds to textual input, we bridge the modality gap between textual and the visual tokens via a learnable MLP model. Next, the projected visual tokens are concatenated with the textual token sequence before being fed to the backbone LLM. Notably, our proposed method is flexible in either integrating with existing VLLM or in being a VLLM on its own. In particular, it is compatible with most of the existing VLLMs architectures, such as LLaMA-VID [30] and VideoLLaMA2 [9]. Here, we denote our method as B-VLLM when there is no integration with other VLLM. In such case, we utilize Qwen2 [56] as backbone LLM for B-VLLM.
@@
[…]
B-VLLM 1fps 50.8 52.1 22.6 51.9 60.8 51.8 47.9 72.3/4.0 59.2/3.5 48.6/3.7
4 Implementation
Baselines. We evaluate against two baselines: LLaMA-VID [30], which uses high spatial compression for video understanding, and VideoLLaMA2 [9], which uniformly samples 8 frames per video. Datasets & Training. B-VLLM is trained solely on LLaMA-VID-Dataset [30] and Valley [38] for fair comparison. See supplementary material for details.